# Local Advice and Local Decompression

### v1

@@ 6.2. $\Delta$-coloring
Hence, $\mathcal{A}$ outputs precisely $\psi$ on the nodes of $G$, which implies that $\mathcal{A}$ outputs a proper $3$-coloring. ∎
8. Structural results and lower bounds
Our main goal here is to show that if we were able to solve all LCLs with some constant number of bits of advice per node, it would violate the Exponential-Time Hypothesis (ETH) (Impagliazzo and Paturi, 1999). However, to do that we need to first show a structural result, which is also of independent interest: algorithms that make use of advice can be made order-invariant, i.e., they do not need the numerical values of the identifiers.
8.1. Order-invariance
In general, in an advice schema our advice bits may depend on the numerical values of the identifiers, and the algorithm $\mathcal{A}$ that solves the problem may make use of the advice bits and the numerical values of the identifiers. However, here we will show that we can essentially for free eliminate the dependency on the numerical values of the identifiers.
A distributed algorithm is called order-invariant (Naor and Stockmeyer, 1995) if the output remains the same if we change the numerical values of the identifiers but preserve their relative order. Put otherwise, an order-invariant algorithm may make use of the graph structure, local inputs (which in our case includes the advice), and the relative order of the identifiers, but not the numerical values of the identifiers.
Naor and Stockmeyer (1995) showed that constant-time distributed algorithms that solve LCL problems can be made order-invariant, and since then order-invariant algorithms have played a key role in many lower-bound results related to constant-time distributed algorithms, see e.g. (Göös et al., 2013, 2017).
We will here prove the following result. We note that it is applicable to LCL problems (they are component-wise defined, and in any LCL we have a bounded $\Delta$ and bounded alphabets $\Sigma_{\mathrm{in}},\Sigma_{\mathrm{out}}$), but also to many other problems that are not locally checkable. We emphasize that we will assume some bound on the maximum degree $\Delta$, but if we have an encoding schema that works for any $\Delta$, we can apply the following result separately for each $\Delta$.
Theorem 1.
Fix a maximum degree $\Delta$. Assume that $\Pi$ is a finite-input component-wise defined graph problem, in which the task is to label nodes with labels from some finite set $\Sigma_{\mathrm{out}}$. Assume that we can solve some graph problem $\Pi$ with some distributed algorithm $\mathcal{A}$ in $T$ rounds using $\beta$ bits of advice. Then we can also solve $\Pi$ with an order-invariant algorithm $\mathcal{A}^{\prime}$ in $T$ rounds using $\beta$ bits of advice.
Proof.
Let $E$ be the encoder that, given a graph $G$ (with some unique identifiers), produces $\beta$-bit advice strings that $\mathcal{A}$ can then use to solve $\Pi$.
In the first steps, we follow the basic idea of Naor and Stockmeyer (1995) to manipulate $\mathcal{A}$. Algorithm $\mathcal{A}$ is a mapping from labeled radius-$T$ neighborhoods to local outputs; we write here $N(v)$ for the radius-$v$ neighborhood of node $v$, and we let $s=\Delta^{T+1}$ be an upper bound on the number of nodes in $N(v)$.
In general, we can encode all information in $N(v)$ as follows:
  1. (1)
    The set of unique identifiers $X\subseteq\{1,2,\dotsc\}$, with $|X(v)|\leq s$.
  2. (2)
So we can reinterpret $\mathcal{A}$ as a function that maps a pair $(X,S)$ to the local output $\mathcal{A}(X,S)\in\Sigma_{\mathrm{out}}$. But we can equally well interpret $\mathcal{A}$ as a function $A$ that maps $X$ to a type $A(X)=F$, where $F$ is a function that maps $S$ to the local output $F(S)\in\Sigma_{\mathrm{out}}$; we simply let $A(X)(S)=\mathcal{A}(X,S)$. The second interpretation turns out to be convenient.
Let us now make our life slightly easier by noticing that we can always pad $X$ with additional identifiers (that use values larger than any value in $X$), so that we will have $|X|=s$.
Then we make the key observation: there are only finitely many different structures $S$. This follows from the assumptions that $\Delta$ is a constant, $\Pi$ is a finite-input problem, and $T$ is a constant. Furthermore, $\Sigma_{\mathrm{out}}$ is finite, so $F$ is a mapping from a finite set to a finite set, and follows that there are only finitely many different types $F$. Let $m$ be the number of possible types, and identify the types with $1,2,\dots,m$.
Now let us stop for a moment and digest what we have learned so far: if our unique identifiers came from the set $I$, then we will have $\mathcal{A}(X,S)=A(X)(S)=F^{*}(S)$. That is, $\mathcal{A}$ will then ignore the numerical values of the identifiers and only pay attention to the structure $S$. However, this is a big if; in general our identifiers can be arbitrary, and even adversarial.
Let us now continue; we will now modify the encoder $E$. We construct a new encoder $E^{\prime}$ that works as follows. Assume we are given a graph $G$, together with some assignment of unique identifiers. The encoder $E^{\prime}$ first constructs graph $G_{1}$ by re-numbering the identifiers (but preserving their order) so that all identifiers in $G_{1}$ come from the monochromatic set $I$.
Now $H$ is a valid instance, and we can feed it to the encoder $E$, which will label it with $\beta$-bit labels so that if we apply $\mathcal{A}$ to $H$ and these labels, we will correctly solve $\Pi$ in each component of $H$. In particular, we will correctly solve $\Pi$ in $G_{1}$. Moreover, in component $G_{1}$, algorithm $\mathcal{A}$ will apply order-invariant algorithm $F^{*}$ in all neighborhoods, ignoring the numerical values of the identifiers.
If we now applied $\mathcal{A}$ to solve $\Pi$ in $G$, it does not necessarily work. However, if we apply $\mathcal{A}^{\prime}(X,S)=F^{*}(S)$ to solve $\Pi$ in $G$, it will behave in exactly the same way as applying $\mathcal{A}$ in $G_{1}$. Hence, $\mathcal{A}^{\prime}$ will also solve $\Pi$ correctly in $G$. Furthermore, $\mathcal{A}^{\prime}$ is by construction order-invariant. ∎
We note that the encoder $E^{\prime}$ constructed above is not practical or efficient (even if the original encoder $E$ is). But as we will see next, merely knowing that $\mathcal{A}^{\prime}$ exists can be sufficient.
8.2. Hardness assuming the Exponential-Time Hypothesis
In Section 4 we saw that in graphs with subexponential growth, any LCL can be solved with one bit of advice per node. We now show that one bit does not suffice in general, assuming the Exponential-Time Hypothesis (Impagliazzo and Paturi, 1999).
Recall that the Exponential-Time Hypothesis states that 3-SAT cannot be solved in time $O(2^{\delta n})$ for some $\delta>0$, where $n$ is the number of variables in the 3-SAT instance. We prove the following:
Theorem 2.
Fix any $\beta$. Assume that for every LCL problem $\Pi$ there is some $T$ such that $\Pi$ can be solved in $T$ rounds with $\beta$ bits of advice. Then the Exponential-Time Hypothesis is false.
Proof.
Consider some 3-SAT instance $\phi$, with $n$ variables and $m$ clauses. By applying the sparsification lemma by (Impagliazzo et al., 2001), we can assume w.l.o.g. that $m=O(n)$.
Now turn $\phi$ into a bipartite graph, with nodes representing variables on one side and nodes representing clauses on the other side. Each node is labeled by its type (variable or clause) and each edge $vc$ is labeled to indicate whether the variable $v$ is negated in the clause $c$. This results in a bipartite graph with $n+m=O(n)$ nodes, and each clause-node has degree at most $3$. Hence, the total number of edges is at most $3m$.
Variable-nodes may have arbitrarily high degrees, but the sum of their degrees is bounded by $3m$. We replace each variable-node that has degree higher than $3$ by a cycle of variable-nodes, with the new edges labeled by equality constraints. This results in a graph in which all nodes have degree at most $3$, and the total number of nodes is bounded by $3m+m=O(n)$. Let $G_{0}$ be the resulting graph, let $n_{0}=O(n)$ be the number of nodes in $G_{0}$, let $\Delta_{0}=3$ be the maximum degree, and let $\ell_{0}=O(1)$ be the number of node and edge labels that we used to encode the instance.
Now assume that we have defined an LCL problem $\Pi_{i}$ and a graph $G_{i}$ with $n_{i}$ nodes, maximum degree $\Delta_{i}$, and $\ell_{i}$ labels; the base case $i=0$ was presented above. We will now define a new LCL problem $\Pi_{i+1}$ and a new graph $G_{i+1}$ as follows. We contract edges in $G_{i}$ to construct $G_{i+1}$ so that we satisfy two properties: each node in $G_{i+1}$ represents $O(1)$ nodes of $G_{i}$, but the number of nodes is $n_{i+1}\leq n_{i}/2$. This can be achieved by e.g. greedily contracting edges, favoring the edges whose endpoints currently represent the smallest number of nodes of $G_{i}$.
We label the nodes of $G_{i+1}$ so that given the input labels of the nodes, we can also recover the original graph $G_{i}$. As each node represents a bounded number of original nodes, and the original graph had maximum degree $\Delta_{i}$, the new graph will have maximum degree $\Delta_{i+1}$ that only depends on $i$. Also, the number of labels $\ell_{i+1}$ will be bounded by a constant that only depends on $i$.
By assumption, there exists a distributed algorithm $\mathcal{A}$ that solves $\Pi_{i}$ with $\beta$ bits of advice per node. Using Theorem 1, we can also assume that $\mathcal{A}$ is order-invariant. Hence, $\mathcal{A}$ is a finite function, and we can compute $\mathcal{A}$ in constant time (where the constant depends on $i$, which depends on $\beta$ and $\varepsilon$, but is independent of $n$).
Now we simply try out all possible strings of advice; there are at most $2^{\varepsilon n}$ such strings. For each advice combination, we try to apply $\mathcal{A}$ to solve $\Pi_{i}$ in $G_{i}$; we simulate $\mathcal{A}$ at each of the $n_{i}=O(n)$ nodes. If and only if $\phi$ is satisfiable, we will find an advice string such that $\mathcal{A}$ succeeds in solving $\Pi_{i}$. The overall running time is $O(2^{\varepsilon n}\cdot n\cdot f(\varepsilon,\beta))$. ∎
9. Composability
@@
Each set $B(v)$ can be encoded by using $c\alpha/\gamma^{3}$ bits, as follows. First, each pair $(i,\ell_{i}(v))$ is encoded by using at most $\lceil\log k\rceil+2\lceil\log(c\alpha/(2k\gamma)^{3})\rceil+1+c\alpha/(2k\gamma)^{3}$ bits by concatenating the bit-string representation of $i$, followed by $\lceil\log|\ell_{i}(v)|\rceil$ $1$s, followed by a 00, followed by the bit-string representation of $|\ell_{i}(v)|$, followed by $\ell_{i}(v)$. Observe that such a string can be decoded to recover the pair. Then, the bit-strings of the pairs are concatenated into a single bit-string $\ell_{G}(v)$, that has length at most $k\lceil\log k\rceil+4kc\alpha/(2k\gamma)^{3}\leq 5kc\alpha/(2k\gamma)^{3}\leq c\alpha/\gamma^{3}$, where the first inequality holds by the definition of $A$. The schema $S$ is defined as the one using, $\ell_{G}(v)$, for each $G\in\mathcal{G}$.
Nodes can solve $\Pi$ as follows. First, each node $v$ decodes its own advice, obtaining a string $\ell_{i}(v)$ for each $1\leq i\leq k$. Then, nodes consider the graph $D$: problems that are sinks in $D$ can be solved in parallel, and after that, nodes can recurse on the subgraph of $D$ induced by problems that still have to be solved. In total, they spend $T(\alpha,\Delta)$ time. ∎
Lemma 0.
@@
Proof.
Let $c$ be a small-enough constant to be fixed later, and let $\alpha=A(c,\gamma)$. Let $S$ be the advice schemas for $\Pi$ obtained from Definition 4 by using parameters $c$ and $\alpha$. Let $G=(V,E)\in\mathcal{G}$. Let $\ell$ be the bit-assignment of the schema $S$ on the graph $G$. We define the schema $S^{\prime}$ by providing a function $\ell^{\prime}$ that maps each node of $V$ into a single bit.
@@
In the following, we define the function $\ell^{\prime}$. At first, define $\ell^{\prime}$ by assigning a $1$ to each bit-holding node, and 00 to all the other nodes. Then, for each cluster $C$, we operate as follows. Let $N(C)$ be the set of nodes at distance at most $d/4$ from the nodes in $C$. Observe that, for two different clusters $C$ and $C^{\prime}$, it holds that the nodes in $N(C)$ are at distance at least $d/2$ from all the nodes in $N(C^{\prime})$. For each node $v\in N(C)$, let $x_{v}$ be the distance between $v$ and its nearest node in $C$. There are two cases:
@@
In the following, we present the bit-assignment for the clusters of the former case. Let $v_{1},\ldots,v_{k}$ be the bit-holding nodes of $C$, sorted by their ID. Recall that the $k\leq\gamma$, and that each bit-holding node has a bit-string of length at most $c\alpha/\gamma^{3}$. Let $B=(L_{1},\ldots,L_{k})$, where $L_{i}=\ell(v_{i})$. We encode the array $B$ by using at most $4c\alpha/\gamma^{2}$ bits, as follows. Each $L_{i}$ is encoded by writing $\lceil\log|L_{i}|\rceil$ $1$s, followed by a 00, followed by the bit-string representation of $|L_{i}|$, followed by $L_{i}$, which requires at most $4c\alpha/\gamma^{3}$ bits. Then, we concatenate the obtained bit-strings, in ascending order of $i$, obtaining a string $L$. Note that, given $L$, it is possible to recover $B$.
Then, we create a new bit-string, by replacing each 00 of $L$ with a sequence of $\gamma+1$ $1$s followed by a 00, and each $1$ of $L$ with a sequence of $\gamma+2$ $1$s followed by a 00. Let $L^{\prime}$ be the obtained bit-string. Note that this string has length at most $(\gamma+3)4c\alpha/\gamma^{2}\leq 16c\alpha/\gamma$, and that it is possible to recover $L$ given $L^{\prime}$.
Recall that $d/8=\alpha/(80\gamma)$. We fix $c$ small enough so that $816c\alpha/\gamma<d/$ 8. Consider the path $P=(v_{1},\ldots,v_{d/8})$ obtained as follows. Node $v_{1}$ is $z$ (that is, the node satisfying $x_{z}\geq d/8+10$), and node $v_{i+1}$ is an arbitrary neighbor of $v_{i}$ satisfying $x_{v_{i+1}}=x_{v_{i}}-1$. We pad the string $L^{\prime}$ to be of length exactly $d/8$ by adding 00s. Then, for each $i$, we set $\ell^{\prime}(v_{i})$ as the $i$th bit of $L^{\prime}$ (possibly overriding the 00 that has been previously assigned).
@@
Finally, observe that, by making $c$ small enough (and possibly increasing $\alpha$ suitably), we can make the ratio between assigned $1$s and assigned 00s an arbitrarily small constant. ∎
Acknowledgements.
This work was initiated at Dagstuhl Seminar 23071: “From Big Data Theory to Big Data Practice”. We would like to thank Sameep Dahal and Augusto Modanese for insightful discussions.

### v2

@@ C.2 $\Delta$-coloring
Hence, $\mathcal{A}$ outputs precisely $\psi$ on the nodes of $G$, which implies that $\mathcal{A}$ outputs a proper $3$-coloring.
Appendix E Composability
@@
Each set $B(v)$ can be encoded by using $c\alpha/\gamma^{3}$ bits, as follows. First, each pair $(i,\ell_{i}(v))$ is encoded by using at most $\lceil\log k\rceil+2\lceil\log(c\alpha/(2k\gamma)^{3})\rceil+1+c\alpha/(2k\gamma)^{3}$ bits by concatenating the bit-string representation of $i$, followed by $\lceil\log|\ell_{i}(v)|\rceil$ 11s, followed by a 0, followed by the bit-string representation of $|\ell_{i}(v)|$, followed by $\ell_{i}(v)$. Observe that such a string can be decoded to recover the pair. Then, the bit-strings of the pairs are concatenated into a single bit-string $\ell_{G}(v)$, that has length at most $k\lceil\log k\rceil+4kc\alpha/(2k\gamma)^{3}\leq 5kc\alpha/(2k\gamma)^{3}\leq c\alpha/\gamma^{3}$, where the first inequality holds by the definition of $A$. The schema $S$ is defined as the one using, $\ell_{G}(v)$, for each $G\in\mathcal{G}$.
Nodes can solve $\Pi$ as follows. First, each node $v$ decodes its own advice, obtaining a string $\ell_{i}(v)$ for each $1\leq i\leq k$. Then, nodes consider the graph $D$: problems that are sinks in $D$ can be solved in parallel, and after that, nodes can recurse on the subgraph of $D$ induced by problems that still have to be solved. In total, they spend $T(\alpha,\Delta)$ time.
Lemma E.3.
@@
Proof E.4.
Let $c$ be a small-enough constant to be fixed later, and let $\alpha=A(c,\gamma)$. Let $S$ be the advice schemas for $\Pi$ obtained from Definition˜3.4 by using parameters $c$ and $\alpha$. Let $G=(V,E)\in\mathcal{G}$. Let $\ell$ be the bit-assignment of the schema $S$ on the graph $G$. We define the schema $S^{\prime}$ by providing a function $\ell^{\prime}$ that maps each node of $V$ into a single bit.
@@
In the following, we define the function $\ell^{\prime}$. At first, define $\ell^{\prime}$ by assigning a $1$ to each bit-holding node, and 0 to all the other nodes. Then, for each cluster $C$, we operate as follows. Let $N(C)$ be the set of nodes at distance at most $d/4$ from the nodes in $C$. Observe that, for two different clusters $C$ and $C^{\prime}$, it holds that the nodes in $N(C)$ are at distance at least $d/2$ from all the nodes in $N(C^{\prime})$. For each node $v\in N(C)$, let $x_{v}$ be the distance between $v$ and its nearest node in $C$. There are two cases:
@@
In the following, we present the bit-assignment for the clusters of the former case. Let $v_{1},\ldots,v_{k}$ be the bit-holding nodes of $C$, sorted by their ID. Recall that the $k\leq\gamma$, and that each bit-holding node has a bit-string of length at most $c\alpha/\gamma^{3}$. Let $B=(L_{1},\ldots,L_{k})$, where $L_{i}=\ell(v_{i})$. We encode the array $B$ by using at most $4c\alpha/\gamma^{2}$ bits, as follows. Each $L_{i}$ is encoded by writing $\lceil\log|L_{i}|\rceil$ 11s, followed by a 0, followed by the bit-string representation of $|L_{i}|$, followed by $L_{i}$, which requires at most $4c\alpha/\gamma^{3}$ bits. Then, we concatenate the obtained bit-strings, in ascending order of $i$, obtaining a string $L$. Note that, given $L$, it is possible to recover $B$.
Then, we create a new bit-string, by replacing each 0 of $L$ with a sequence of $\gamma+1$ 11s followed by a 0, and each $1$ of $L$ with a sequence of $\gamma+2$ 11s followed by a 0. Let $L^{\prime}$ be the obtained bit-string. Note that this string has length at most $(\gamma+3)4c\alpha/\gamma^{2}\leq 16c\alpha/\gamma$, and that it is possible to recover $L$ given $L^{\prime}$.
Recall that $d/8=\alpha/(80\gamma)$. We fix $c$ small enough so that $16c\alpha/\gamma<d/8$. Consider the path $P=(v_{1},\ldots,v_{d/8})$ obtained as follows. Node $v_{1}$ is $z$ (that is, the node satisfying $x_{z}\geq d/8+10$), and node $v_{i+1}$ is an arbitrary neighbor of $v_{i}$ satisfying $x_{v_{i+1}}=x_{v_{i}}-1$. We pad the string $L^{\prime}$ to be of length exactly $d/8$ by adding 0s. Then, for each $i$, we set $\ell^{\prime}(v_{i})$ as the $i$th bit of $L^{\prime}$ (possibly overriding the 0 that has been previously assigned).
@@
Finally, observe that, by making $c$ small enough (and possibly increasing $\alpha$ suitably), we can make the ratio between assigned $1$s and assigned 0s an arbitrarily small constant.